# GAPS: Guiding Dynamic Android Analysis with Static Path Synthesis

### v1

@@ 5.4. Backward Generation of the Call Graph
GAPS adopts a goal-directed, demand-driven approach to call graph construction that departs from traditional whole-program analyses, such as the one used by Flowdroid. Instead of building a complete call graph upfront, GAPS performs a targeted, backward traversal starting from a specified method of interest. This traversal is both inter-procedural and context-sensitive, collecting only those method calls that lie along feasible paths leading to the target. The resulting structure is a partial call graph slice that is sufficient for reachability reasoning but avoids the scalability bottlenecks of global analyses. GAPS also incorporates data-flow analysis to capture branch conditions and control dependencies essential for reconstructing realistic execution paths, as well as actionable entry points and GUI event identifiers. The pseudocode for the partial call graph generation can be consulted in Algorithm 2.
The process starts from the target method and iteratively traverses backwards through its potential callers. This is implemented using a queue of nodes to process (2 and 3), where the current node being processed is removed from a queue (6) and all its statically discoverable callers are identified via find_new_nodes (10). These new nodes are added to the call graph and appended to the queue (11). When an entry point is encountered, it serves to terminate the backward traversal, and no further predecessor nodes are added. Then, nodes that are processed are added to the seen_nodes set (12), to avoid reprocessing and ensure termination (7). This backward expansion continues until no new nodes are found. At that point, the constructed call graph contains a slice of the original application call graph that is specifically focused on reaching the target. Once the partial call graph is complete, GAPS invokes a depth-first traversal (dfs_visit) (14) to extract all feasible paths from known application entry points to the target method. These paths are further enriched by resolving conditional branches (16 and 17) and detecting associated GUI elements (18), which are necessary to trigger the GUI events.
@@ 5.5. Linking paths that satisfy conditional statements
The illustrative example contains a conditional statement involving the access variable that will be analyzed to find the paths that set its value to True. This happens by contacting the AccessReceiver component. A separate path will be reconstructed and prepended to the main one to reproduce this.
@@ 5.6. GUI Event Retrieval
To support guided dynamic execution, GAPS must identify the specific GUI elements responsible for triggering event handlers within the app. This enables the generation of interaction plans that simulate user actions accurately. The retrieval of these GUI elements is performed statically and relies on a layered application of points-to analysis. Algorithm 4 illustrates the procedure.
The process begins by detecting GUI event handlers, such as those associated with callbacks like onClick() or onItemSelected(), that happen in a recognized activity. GAPS scans the paths for listener registrations, including explicit calls to methods like setOnClickListener(). Once a handler is found, points-to analysis is used to resolve the Android object instance linked to the callback (2), such as a Button, MenuItem, or other view element. After identifying the GUI object, the paths where it is involved are retrieved (3), and an iterative round of points-to analysis is used to trace its instantiation back to common methods such as findViewById(). From this trace, GAPS extracts the associated resource identifier (e.g., button_main) by analyzing the arguments passed to the call (5). The final outcome of this analysis is a list mapping the activity and the view identifier, representing a GUI event. These events are later embedded into high-level interaction instructions, allowing the dynamic executor to match statically derived GUI elements with actual runtime views.
@@
10   gui_bindings.append(activty, gui_id)
@@ 6. Implementation
In this section, we provide more details about GAPS’ implementation: the call graph generation and path reconstruction, Graphical Elements’ ID retrieval, and the automatic interaction module.
@@ 6.1. Call Graph Generation and Path Reconstruction
At its core, GAPS uses the Androguard (androguard) Python library to load and analyze the APK or DEX files in input through the AnalyzeAPK and AnalyzeDex (analyzeapk) built-in functions. Then, GAPS analyzes the MethodAnalysis construct for each method to access its basic blocks. Every basic block obtained is represented by a sequence of smali instructions, an intermediate representation of the app bytecode. Internally, GAPS uses the networkx (networkx) Python library to create and model the graph. Subsequently, we use the all_shortest_paths (all_shortest_paths) method from the networkx library to retrieve the paths connecting the source nodes and the entry points. Since solving implicit flows is still considered an open research problem (cg_soundness) that falls out of the scope of this paper, we have added callback mappings provided by popular works in Android static analyzers, such as EdgeMiner (edgeminer) and Soot (soot) through its virtual edges (virtualedges). In this way, we can find the corresponding constructor or method initializer and partially solve implicit flows. Data-flow analysis is performed following the smali code representation, which provides a rigorous register-based syntax that we modeled to adapt the points-to-analysis and constant propagation. For example, to track the transition to a new activity, when an Intent is constructed, the last argument passed to the constructor corresponds to the destination activity, which matches the last register in smali syntax. This data-flow-aware analysis is also essential for linking paths that satisfy conditional statements. Moreover, through constant propagation, a conditional statement involving primitive-type variables, objects, or method return values is translated to compare constant values to test its satisfiability.
@@ 6.2. Static Output and Dynamic Guidance
GAPS uses Apktool (apktool) to decompile apps and retrieve GUI resource IDs during static analysis. To convert them from the smali numerical representation, GAPS consults a file generated by Apktool called public.xml, where they can be converted into a literal one. GAPS can generate high-level instructions in JSON files, consisting of actionable entry points and sequences of graphical element IDs to interact with, along with the Activity name where they reside. Additionally, a call sequence is also included. These results can be used in two ways. First, they can serve as the standalone output of the static analysis phase: for each target method, GAPS lists all the paths found during analysis, together with their reconstructed call sequence and the corresponding high-level interaction plan. Second, they can directly guide dynamic analysis. In this case, each plan begins with instructions that can be executed through adb to send an Intent and trigger the correct entry point discovered during the static analysis. The plan then continues with a sequence of UI element IDs and the Activity in which they appear, specifying the concrete actions to perform on the app interface in order to reach the target method. When conditional statements are resolved during the static phase, the relative call sequences and their associated high-level instructions are prepended to the plan, ensuring that the conditions are satisfied before continuing with the UI interactions.
@@ 6.3. Automatic Interaction
The high-level instructions generated through the static analysis allow for dynamic traversal of the paths by (i) using one of the app’s entry points and (ii) graphically interacting with the GUI. More specifically, the GAPS dynamic module relies on the AndroidViewClient (androidviewclient) Python library to find the graphical elements through the findViewById (avc1) and touch (avc2) APIs, given their literal identifiers, and subsequently interact with them. Additionally, Frida (frida), a dynamic instrumentation framework, can be used to apply a hook on the target method and get a confirmation when it is reached.
The previous approach comes with some limitations for apps that display unpredictable elements in an overlay on the screen. For example, many apps can show dynamically rendered content or require the supply of textual input. To address the unpredictability of some apps’ UI components, we integrated Guardian (guardian) in the GAPS’ automatic interaction module: when GAPS cannot locate an activity or widget on the screen, Guardian guides the interaction with the app until both the widget and the activity appear, enabling GAPS to resume its automatic interaction under path guidance. This choice is motivated by the comparison in subsection 4.2, where Guardian performed better than APE and GoalExplorer thanks to its LLM capabilities, which enable goal-driven interaction planning and more effective UI navigation, including the generation of textual input. For instance, applications that require semantic planning, such as Yahtzee, where the number of rounds and players must be specified in the very first activity, are better suited for exploration through LLM-based UI interaction tools like Guardian. We include Guardian as an optional component, which can be disabled or replaced with more advanced tools as they become available in the future. Furthermore, GAPS uses memorization to save the sequence of GUI events performed by Guardian when it is able to transition between two activities, allowing GAPS to reproduce the steps independently.
@@ 7. Results
In this section, we present the results of our experimental evaluation of GAPS. In subsection 7.1, we compare GAPS against representative static and dynamic tools using the AndroTest benchmark, evaluating both its static path synthesis and dynamic guided execution capabilities. Finally, in subsection 7.2 we present a case study on 25 real-world applications attempting to reach security-relevant code. The target methods, in this case, are provided by the SPECK (speck) rule-based static analyzer.
7.1. RQ3: Effectiveness of GAPS Compared to Existing Tools
To evaluate whether GAPS improves the reachability of target methods over state-of-the-art tools, we compare it against two static analyzers (FlowDroid and DroidReach) and three dynamic GUI testers (GoalExplorer, APE, and Guardian). GAPS is evaluated in both static and dynamic configurations: the static phase measures the percentage of methods for which GAPS can generate paths, while the dynamic phase measures actual runtime reachability during guided execution.
Table 1. Comparison of target-method reachability and analysis time across static and dynamic tools.
@@ 7.1. RQ3: Effectiveness of GAPS Compared to Existing Tools
Table 1 presents a comparative analysis of target-method reachability and average analysis time for static and dynamic tools. GAPS significantly outperforms both classes of tools. Its static pre-processing achieves 88.24% reachability by generating partial paths that satisfy known conditions, entry points, and provides GUI events. Furthermore, its analysis time, on average, is 81.8% and 87.8% faster compared to DroidReach and FlowDroid, respectively. Meanwhile, its dynamic component, guided by statically synthesized interaction plans, reaches 57.44% of targets during execution, substantially higher than both APE and Guardian.
On average, Guardian was invoked four times per app, which corresponds to contributing to reaching 9.3% of the methods, indicating that GAPS primarily relies on its static analysis module for guidance, with Guardian used only as a fallback.
RQ3: Effectiveness of GAPS Compared to Existing Tools GAPS outperforms static and dynamic state-of-the-art tools both in terms of analysis time and target-method reachability. Its static phase covers 88.24% of target methods with an average analysis time of 4.27 seconds, while dynamic execution reaches 57.44%.
7.2. RQ4: GAPS’ Applicability to Real-World Use Cases
To demonstrate the practical utility of GAPS, we evaluated its performance on the 50 most downloaded real-world applications from the Google Play Store, collected in August 2025. The full list of applications and the results can be found in our repository [^3] https://github.com/samudoria/GAPS/tree/main/evaluation_data/real_world_apps_results/. It should be noted that, in this case, AndroLogs (samhi2024androlog) failed to repackage the real-world apps to add code coverage feedback. Instead, we used GAPS’ Frida integration to hook the target methods and log whenever they were reached.
GAPS achieved a static reachability of 62.03% of the target methods, with an average analysis time of 278.9 seconds (less than 5 minutes) per app. This performance highlights GAPS’s ability to scale to large, complex applications while maintaining efficiency. The lower reachability compared to the AndroTest benchmark can be attributed to the increased complexity and obfuscation techniques employed in commercial apps, as well as the use of third-party frameworks which obscure traditional control-flow patterns. During dynamic execution, GAPS successfully reached 59.86% of the target methods. The gap between static and dynamic reachability arises from challenges such as when a complex app state is required (e.g., performing an app payment), dynamic UI rendering, and the need for specific input data.
RQ4: GAPS’ Applicability to Real-World Use Cases GAPS was evaluated on 50 top Google Play Store apps, demonstrating practical utility and scalability in real-world scenarios. The tool achieved 62.03% static reachability of target methods with an average analysis time of 278.9 seconds (less than 5 minutes per app), and 59.86% dynamic reachability.
@@ 7.3. Discussion and Limitations
First, its reliance on static bytecode analysis restricts its effectiveness for apps built with frameworks like Flutter or React Native, where the app’s logic resides outside traditional Dalvik code. Furthermore, apps with an obfuscated call graph can lead to path explosion and significantly slow down analysis time, which can be noted from the higher analysis time of real-world apps. Another obstacle to path reconstruction is dead code located in libraries, which can worsen GAPS’ performance. Finally, call graph generation is known to suffer from unsoundness (cg_soundness), caused by untracked implicit flows. GAPS uses the same callback mappings used by FlowDroid (flowdroid) and Soot (soot), and also includes those generated by EdgeMiner (edgeminer), but might still suffer from unsoundness when an unexpected implicit flow is met. The analysis of the Android Framework to mine implicit edges for static analysis tools is an orthogonal problem that is out of scope for this work. Our current approach does not support Jetpack Compose. Unlike XML-based layouts, Jetpack Compose defines UI elements directly in Kotlin code and attaches event handlers through lambdas rather than OnClickListener objects. Furthermore, UI state can change dynamically through recomposition. These differences make data-flow analysis for Jetpack Compose more complex. Extending static analysis to reliably extract GUI elements and events from Jetpack Compose remains an open research problem, and we consider this an important direction for future work.
In some cases, we cannot dynamically attest the reachability. Complex interactions might be required to fully traverse a path: Hot Death and Yahtzee are games where some of the randomly selected target methods can only be reached upon winning a game. Finally, most apps have paths with intent filters as an entry point, which can have permissions or be owned by the system (e.g., BOOT_COMPLETED), or could require data to be passed along with the Intent. Retrieving the required data parameters and automatically generating Intents is an orthogonal problem to dynamic analysis that falls out of the scope of this paper.
Despite these challenges, GAPS’s hybrid approach balances precision and scalability, offering a practical foundation for future enhancements.

### v2

@@ 5.4. Backward Generation of the Call Graph
GAPS adopts a goal-directed, demand-driven approach to call graph construction that diverges from traditional whole-program analyses, such as those used by Flowdroid. Instead of building a complete call graph upfront, GAPS performs a targeted, backward traversal starting from a specified method of interest. This traversal is both inter-procedural and context-sensitive, collecting only those method calls that lie along feasible paths leading to the target. The resulting structure is a partial call graph slice that is sufficient for reachability purposes but avoids the scalability bottlenecks of global analyses. GAPS also incorporates data-flow analysis to capture branch conditions and control dependencies, which are essential for reconstructing realistic execution paths, as well as actionable entry points and GUI event identifiers. The pseudocode for the partial call graph generation can be consulted in Algorithm 2.
The process begins with the target method and iteratively traverses backward through its potential callers. This is implemented using a queue of nodes to process (2 and 3), where the current node being processed is removed from a queue (6) and all its statically discoverable callers are identified via find_new_nodes (10). These new nodes are added to the call graph and appended to the queue (11). When an entry point is encountered, it serves to terminate the backward traversal, and no further predecessor nodes are added. Then, nodes that are processed are added to the seen_nodes set (12), to avoid reprocessing and ensure termination (7). This backward expansion continues until no new nodes are found. At that point, the constructed call graph contains a slice of the original application call graph that is specifically focused on reaching the target. Once the partial call graph is complete, GAPS invokes a depth-first traversal (dfs_visit) (14) to extract all feasible paths from known application entry points to the target method. These paths are further enriched by resolving conditional branches (16 and 17) and detecting associated GUI elements (18), which are necessary to trigger the GUI events.
@@ 5.5. Linking paths that satisfy conditional statements
The illustrative example contains a conditional statement involving the access variable, which will be analyzed to identify the paths that set its value to true. This happens by contacting the AccessReceiver component. A separate path will be reconstructed and prepended to the main one to reproduce this.
@@ 5.6. GUI Event Retrieval
To support guided dynamic execution, GAPS must identify the specific GUI elements responsible for triggering event handlers within the app. This enables the generation of interaction plans that accurately simulate user actions. The retrieval of these GUI elements is performed statically and relies on a layered application of points-to analysis. Algorithm 4 illustrates the procedure.
The process begins by detecting GUI event handlers, such as those associated with callbacks like onClick() or onItemSelected(), that happen in a recognized activity. GAPS scans the paths for listener registrations, including explicit calls to methods like setOnClickListener(). Once a handler is found, points-to analysis is used to resolve the Android object instance linked to the callback (2), such as a Button, MenuItem, or other view element. After identifying the GUI object, the paths in which it is involved are retrieved (3), and an iterative round of points-to analysis is used to trace its instantiation back to common methods, such as findViewById(). From this trace, GAPS extracts the associated resource identifier (e.g., button_main) by analyzing the arguments passed to the call (5). The final outcome of this analysis is a list mapping the activity and the view identifier, representing a GUI event. These events are later embedded into high-level interaction instructions, allowing the dynamic executor to match statically derived GUI elements with actual runtime views.
@@
10   gui_bindings.append(activity, gui_id)
@@ 6. Implementation
In this section, we provide more details about GAPS’ implementation, including call graph generation and path reconstruction, retrieval of Graphical Elements’ IDs, and the automatic interaction module.
@@ 6.1. Call Graph Generation and Path Reconstruction
At its core, GAPS uses the Androguard (Androguard, 2025b) Python library to load and analyze the APK or DEX files in input through the AnalyzeAPK and AnalyzeDex (Androguard, 2025a) built-in functions. Then, GAPS analyzes the MethodAnalysis construct for each method to access its basic blocks. Every basic block obtained is represented by a sequence of smali instructions, an intermediate representation of the app bytecode. Internally, GAPS uses the networkx (NetworkX, 2025b) Python library to create and model the graph. Subsequently, we use the all_shortest_paths (NetworkX, 2025a) method from the networkx library to retrieve the paths connecting the source nodes and the entry points. Since solving implicit flows is still considered an open research problem (Samhi et al., 2024) that falls out of the scope of this paper, we have added callback mappings provided by popular works in Android static analyzers, such as EdgeMiner (Cao et al., 2015) and Soot (Soot, 2025b) through its virtual edges (Soot, 2025a). In this way, we can find the corresponding constructor or method initializer and partially solve implicit flows. Data-flow analysis is performed following the smali code representation, which provides a rigorous register-based syntax that we have modeled to adapt the points-to analysis and constant propagation. For example, to track the transition to a new activity, when an Intent is constructed, the last argument passed to the constructor corresponds to the destination activity, which matches the last register in smali syntax. This data-flow-aware analysis is also essential for linking paths that satisfy conditional statements. Moreover, through constant propagation, a conditional statement involving primitive-type variables, objects, or method return values is translated to compare constant values to test its satisfiability.
@@ 6.2. Static Output and Dynamic Guidance
GAPS uses Apktool (APKTool, 2025) to decompile apps and retrieve GUI resource IDs during static analysis. To convert them from the smali numerical representation, GAPS consults a file generated by Apktool called public.xml, where they can be converted into a literal one. GAPS can generate high-level instructions in JSON files, consisting of actionable entry points and sequences of graphical element IDs to interact with, along with the Activity name where they reside. Additionally, a call sequence is also included. These results can be used in two ways. First, they can serve as the standalone output of the static analysis phase: for each target method, GAPS lists all the paths found during analysis, together with their reconstructed call sequence and the corresponding high-level interaction plan. Second, they can directly guide dynamic analysis. In this case, each plan begins with instructions that can be executed through adb to send an Intent and trigger the correct entry point discovered during the static analysis. The plan then continues with a sequence of UI element IDs and the Activity in which they appear, specifying the concrete actions to perform on the app interface in order to reach the target method. When conditional statements are resolved during the static phase, the relative call sequences and their associated high-level instructions are prepended to the plan, ensuring that the conditions are satisfied before continuing with the UI interactions.
@@ 6.3. Automatic Interaction
The high-level instructions generated through the static analysis allow for dynamic traversal of the paths by (i) using one of the app’s entry points and (ii) graphically interacting with the GUI. More specifically, the GAPS dynamic module relies on the AndroidViewClient (dtmilano, 2025c) Python library to find the graphical elements through the findViewById (dtmilano, 2025a) and touch (dtmilano, 2025b) APIs, given their literal identifiers, and subsequently interact with them. Additionally, Frida (Frida, 2025), a dynamic instrumentation framework, can be used to apply a hook on the target method and get a confirmation when it is executed.
The previous approach comes with some limitations for apps that display unpredictable elements in an overlay on the screen. For example, many apps can display dynamically rendered content or require the input of textual information. To address the unpredictability of some apps’ UI components, we integrated Guardian (Ran et al., 2024) in the GAPS’ automatic interaction module: when GAPS cannot locate an activity or widget on the screen, Guardian guides the interaction with the app until both the widget and the activity appear, enabling GAPS to resume its automatic interaction under path guidance. This choice is motivated by the comparison in subsection 4.2, where Guardian outperformed APE and GoalExplorer, thanks to its LLM capabilities, which enable goal-driven interaction planning and more effective UI navigation, including the generation of textual input. For instance, applications that require semantic planning, such as Yahtzee, where the number of rounds and players must be specified in the very first activity, are better suited for exploration through LLM-based UI interaction tools like Guardian. We include Guardian as an optional component, which can be disabled or replaced with more advanced tools as they become available in the future. Furthermore, GAPS uses memorization to save the sequence of GUI events performed by Guardian when it is able to transition between two activities, allowing GAPS to reproduce the steps independently.
@@ 7. Results
In this section, we present the results of our experimental evaluation of GAPS. In subsection 7.1, we compare GAPS against representative static and dynamic tools using the AndroTest benchmark, evaluating both its static path synthesis and dynamic guided execution capabilities. Finally, in subsection 7.2, we present a case study on 50 real-world apps, by using GAPS to execute possible vulnerable methods identified through SPECK (SPRITZ, 2025), a rule-based static analyzer.
7.1. RQ3: Comparison of GAPS with State-of-Art Tools
To evaluate GAPS’ capability to statically generate paths towards target methods and to dynamically execute them, we compare it against two static analyzers (FlowDroid and DroidReach) and three dynamic GUI testers (GoalExplorer, APE, and Guardian). Our evaluation involves both static and dynamic configurations: the static one measures the percentage of methods for which GAPS can generate paths, while the dynamic one measures actual runtime method execution.
Table 1. Comparison of GAPS with static and dynamic analysis tools. For static tools, the reachability refers to the number of target methods for which a path has been statically generated. For dynamic tools, the reachability refers to the number of target methods executed at runtime.
@@ 7.1. RQ3: Comparison of GAPS with State-of-Art Tools
Table 1 presents a comparative analysis, in terms of target-method reachability and average analysis time, for static and dynamic tools. GAPS significantly outperforms both classes of tools. Its static module is able to generate paths towards 88.24% of methods. Furthermore, GAPS’ static analysis time, on average, is 81.8% and 87.8% faster compared to DroidReach and FlowDroid, respectively. Meanwhile, its dynamic module, guided by the statically precomputed paths, can execute 57.44% of target methods, substantially outperforming both APE and Guardian.
On average, Guardian was invoked four times per app, contributing to reaching 9.3% of the methods. Thus, this confirms that GAPS primarily relies on its static analysis module for guidance, with Guardian used only as a fallback.
RQ3: Comparison of GAPS with State-of-Art Tools GAPS outperforms static and dynamic state-of-the-art tools both in terms of analysis time and target-method reachability. Its static phase generates paths towards 88.24% of target methods with an average analysis time of 4.27 seconds, while the dynamic execution triggers 57.44% of the methods.
7.2. RQ4: GAPS’ Usage on Real-World Apps
To evaluate GAPS on real-world apps, we launched it over the 50 most downloaded apps from the Google Play Store, collected in August 2025 (the full list of apps and the results can be found in our repository [^3] https://github.com/samudoria/GAPS/tree/main/evaluation_data/real_world_apps_results). It should be noted that, in this case, AndroLogs (Samhi and Zeller, 2024) failed to repackage the real-world apps to add code coverage feedback. Alternatively, we used GAPS’ Frida integration to hook the target methods and log whenever they were executed.
GAPS generated paths towards 62.03% of the target methods, with an average analysis time of 278.9 seconds (less than 5 minutes) per app. This performance highlights GAPS ability to scale to large, complex applications while maintaining efficiency. The lower performance compared to the one achieved on the AndroTest benchmark can be attributed to the increased complexity and obfuscation techniques employed in commercial apps, as well as the use of third-party frameworks, which obscure traditional control-flow patterns. During dynamic analysis, GAPS successfully executed 59.86% of the target methods. The performance gap between the static and dynamic modules arises from challenges such as reconstructing a complex app state (e.g., performing an app payment), dynamic UI rendering, and need for specific input data.
RQ4: GAPS’ Usage on Real-World Apps GAPS was evaluated on 50 top Google Play Store apps, demonstrating practical utility and scalability in real-world scenarios. The tool generated paths towards 62.03% of the target methods with an average analysis time of 278.9 seconds (less than 5 minutes per app), and dynamically executed 59.86% of them.
@@ 7.3. Discussion and Limitations
First, its reliance on static bytecode analysis restricts its effectiveness for apps built with frameworks like Flutter or React Native, where the app’s logic resides outside traditional Dalvik code. Furthermore, apps with an obfuscated call graph can lead to path explosion and significantly slow down analysis time, as evidenced by the higher analysis times of real-world apps. Another obstacle to path reconstruction is dead code located in libraries, which can worsen GAPS’ performance. Finally, call graph generation is known to suffer from unsoundness (Samhi et al., 2024), caused by untracked implicit flows. GAPS uses the same callback mappings used by FlowDroid (Arzt et al., 2014) and Soot (Soot, 2025b), and also includes those generated by EdgeMiner (Cao et al., 2015), but might still suffer from unsoundness when an unexpected implicit flow is met. The analysis of the Android Framework to identify implicit edges for static analysis tools is an orthogonal problem that falls outside the scope of this work. Our current approach does not support Jetpack Compose. Unlike XML-based layouts, Jetpack Compose defines UI elements directly in Kotlin code and attaches event handlers through lambdas rather than OnClickListener objects. Furthermore, UI state can change dynamically through recomposition. These differences make data-flow analysis for Jetpack Compose more complex. Extending static analysis to reliably extract GUI elements and events from Jetpack Compose remains an open research problem, and we consider this an important direction for future work.
In some cases, we cannot dynamically execute target methods, even though GAPS has built a static path. Complex interactions might be required to fully traverse a path: Hot Death and Yahtzee are games where some of the randomly selected target methods can only be reached upon winning a game. Finally, most apps have paths with intent filters as an entry point, which can have permissions or be owned by the system (e.g., BOOT_COMPLETED), or could require data to be passed along with the Intent. Retrieving the required data parameters and automatically generating Intents is an orthogonal problem to dynamic analysis, which falls outside the scope of this paper.
Despite these challenges, GAPS’ hybrid approach strikes a balance between precision and scalability, providing a practical foundation for future enhancements.